The terminal voltage of a 2-H inductor is v = 10(1−t) V. Find the current
flowing through it at t = 4 s and the energy stored in it within 0
4 s. Assume i(0) = 2 A.

Grade:11

2 Answers

raju
8 years ago
(a) Under dc conditions, we replace the capacitor with an open circuit
and the inductor with a short circuit,It is evident from
that
i = i L =i1 Ω5 Ωi L12 V4 Ω+−12= 2A1 + 52 H+v C−1 F
(a) i The voltage v C is the same as the voltage across the 5- resistor. Hence,
1 Ω 5 Ωi Lv C = 5i = 10 V
(b) The energy in the capacitor is
w C =12 V4 Ω+−+v C−1 21Cv = (1)(10 2 ) = 50 J2 C2
(b) and that in the inductor is
w L =1 21Li = (2)(2 2 ) = 4 J
Junel Albao
2 years ago

The terminal voltage of a 2- H inductor is v = 10(1-t) v . Find the current flowing through it at t= 4s and the energy atored in it at t= 4 S. Assume I(0) =2A. Please provide the detailed solution?